A 53 N force is needed to slide a 50.0kg box across a flate surface at a constant velocity. What is the coefficient of kinetic friction between the box and the floor?

0.10, 0.13, 0.09, 0.11
Answer #1


Since the box is moving with constant velocity...means that the body is in dynamic equilibrium ..net force acting on box =0 N

kinetic friction force must equal to 53 N but opposite in direction

then kinetic frictional force = 53 N


mu_k*m*g = 53

coefficient of friction is mu_k = 53/(m*g) = 53/(50*9.8) = 0.108